# Scaling Efficient LLMs

### v4

@@ 5 Recurrent Transformers
With our theoretical results as backdrop, we propose an architecture that converts any transformer into a recurrent network in order to combine the efficacy of the attention mechanism in the transformer, with the efficiency of recurrent networks. Specifically, let $X=(x_{1},...x_{i},...x_{N})$ be an input sequence, where each $x_{i}$ is a block of vectors embedding the corresponding sequence of non-overlapping and contiguous input tokens $w_{i,1},w_{i,2},...w_{i,K}$, with $K\geq 1$ being the block size. Let $\tau$ denote the function computed by a regular transformer, e.g., Vaswani et al. (2017), in that on input sequence $X$ the transformer outputs $\tau(X)=(y_{1},y_{2},...y_{N})$. We assume that the $x_{i}$ and $y_{i}$ consist of vectors of the same dimension, and that $\tau$ preserves causality in that $y_{i}$ does not depend on $x_{i+1},x_{i+2}...x_{N}$. The recurrent transformer based on $\tau$ is as follows, where the comma between two sequences concatenates them:
@@ 5.1 Computational Complexity
Let $L,D$ and $N$ be the number of layers, the embedding dimension, and the sequence length respectively of a regular transformer and $K$ the block size of a recurrent transformer. We note that the computational complexity of the regular transformer is $L(N^{2}D+ND^{2})$, from which it is easy to see that of the recurrent transformer is $L(4DNK+2ND^{2})$.
@@ 6 Experimental Results
We now test the performance of recurrent transformers on several datasets. The goal of our experiments is to test whether smaller single-layer recurrent transformers can match the performance of larger multi-layer transformers at lower computational cost. Owing to budge constraints, all of our experiments were run on an M4 Mac Mini with 16GB memory.
@@ 6.1 Long Range Image Classification
We compare the performance of the recurrent transformer against regular transformers on the “Long Range Arena” image classification dataset of Tay et al. (2020):
“This task is an image classification task, where the inputs are sequences of pixels. In other words, an N × N image is flattened to a sequence of length N2 pixels. …To simplify the setup, we map the input images to a single gray-scale channel where each pixel is represented with an 8-bit pixel intensity (vocabulary size of 256). In LRA, we use the CIFAR-10 dataset…”
Specifically, the CIFAR-10 dataset, Krizhevsky et al. (2009), comprises a training set of 50,000 32x32 pixel images, and 10,000 test images across 10 categories. For this task, the images are converted into 256 grayscale values and fed as an input sequence of 1024 tokens across a vocabulary of 256. The output is a single token across a vocabulary of 10 items representing the classification label of the images. Table 1 specifies the base transformer layer in our comparison.
@@
Regular Transformer 4 40 91K 1.38E8 1.68 (0.02) 41.2 (0.7)
Regular Transformer 1 40 53K 3.46E7 1.75 (0.02) 39.3 (0.8)
Recurrent Transformer 1 120 53K 6.29E6 1.50 (0.01) 47.3 (0.4)
@@
Table 2 shows the results of our experiments for this dataset, comparing a 4-layer and 1-layer regular transformer against a recurrent transformer with block size $K=32$. The "Ops/sample" column reflects our computational complexity estimates of section 5.1. The Validation Loss reported is the average across 100 samples and in parentheses, the corresponding standard error, e.g., 1.5 (0.01) implies that with 95% confidence, the validation loss is within $\pm$ 0.02 of 1.5. All three models were run for 600,000 samples, and the recurrent transformer could handle larger batch sizes than the 4-layer transformer in the available memory. The recurrent transformer outperforms on loss and accuracy, requiring less than $5\%$ of the compute of the 4-layer regular transformer. Of interest, the parameter $\alpha$ of Equation 7 converges rapidly to $\approx 0.45$ during training, suggesting accumulation of history aids in solving this task. Figure 1 shows the loss and accuracy during training.
@@ 6.2 Copy and Selective Copy with Curriculum Training
We test the performance of the recurrent transformer on the copy and selective copy tasks of Gu & Dao (2023). The copy task involves 16 tokens, followed by 4096 noise tokens, followed by 16 tokens to trigger recall of the first 16 tokens, for a total sequence of 4128 tokens. The tokens are drawn from an alphabet of 16 unique characters, of which 2 characters are reserved for the noise and recall tokens respectively. Successful copy would reproduce the first 16 tokens in the original order against the last 16 recall tokens. The selective copy task randomly intersperses the 16 tokens to be copied amongst 4096 noise tokens, followed by 16 tokens to trigger recall of the 16 interspersed tokens, for a total sequence of 4128 tokens.. Successful copy would reproduce the 16 interspersed tokens in their original order against the last 16 recall tokens.
It is known that copy tasks are easy for transformers, but challenging for recurrent networks, requiring $\approx 100M$ parameters, e.g., Gu & Dao (2023), Ren et al. (2024). In our experiments, the surprisingly small recurrent transformer of Table 3 with 0.5M parameters solves the tasks, with block size $K=32$.
@@
Training the recurrent transformer on the copy tasks is challenging due to vanishing gradients, e.g., Hochreiter (1991). We overcome this challenge with curriculum training, e.g., Sodhani et al. (2020). Specifically, the training process begins with say, 128 noise tokens, trains until the validation loss drops down to say 0.3, then doubles the number of noise tokens. This process is repeated until the number of noise tokens reaches 4096, at which point the training process drives the validation loss down to the desired minimal level for completion. During training, the learning rate decays as a function of the sequence length and the validation loss.
Table 4 shows the results of our experiments. The Os/sample reported in the table is when each sample has 4096 noise tokens, which is the case during inference. The Validation Loss and Validation Accuracy are averaged over 100 samples and the numbers in parentheses are the respective standard error. For comparison, the estimated Ops/sample is $\approx 1\%$ of a regular transformer with four layers of Table 3. The values of the accumulation parameter $\alpha$ in the table suggest a balance between forget and accumulate to solve this task. Figure 2 shows the loss and accuracy for the selective copy task during the training process. The spikes in the figures are the result of the shocks introduced by doubling the number of noise tokens during curriculum training. The validation loss and accuracy are better than the corresponding training loss and accuracy due to the dropout during training.
@@ 6.3 Natural Language Processing
We compare the performance of the recurrent transformer against regular transformers on the nanoGPT Shakespeare model and dataset of Karpathy (2023). Table 5 specifies the base transformer layer in our comparison. The recurrent transformer has a block size $K=16$. Table 6 shows the results of our experiments comparing the performance of a 1-layer and 6-layer regular transformer operating on the full sequence length, against a 1-layer recurrent transformer. The Validation Loss is averaged over 40 samples and the number in parentheses is the respective standard error. The Ops/sample entries reflect the computational complexity estimates of Section 5.1. Figure 3 shows the Training Loss and Validation Loss for our experiments against the total training cost per Section 5.1. Although a fifth of the size, the recurrent transformer matches the performance of the multi-layer regular transformer in this experiment, requiring more samples to train but at a lower cost per sample, and therefore about the same total training cost. However, the recurrent transformer requires only about $20\%$ the inference cost of the multi-layer regular transformer. Of interest, the parameter $\alpha$ of Equation 7 converges rapidly to zero during training. Owing to budge constraints, we are unable to test larger models.
@@
Regular Transformer 6 64 11M 3.8E8 1.47 (0.01)
Regular Transformer 1 64 1.9M 6.3E7 1.58 (0.01)
Recurrent Transformer 1 64 1.9M 8.2E7 1.47 (0.01)
@@ 7 Summary
Trained LLMs in the transformer architecture are typically sparse in that most of the parameters are negligible, raising questions on efficiency. Furthermore, the so-called "AI scaling law" for transformers suggests that the number of parameters must scale linearly with the size of the data. In response, we inquire into efficient LLMs, i.e. those with the fewest parameters that achieve the desired accuracy on a training corpus. Specifically, by comparing theoretical and empirical estimates of the Kullback-Liebler divergence, we derive a natural AI scaling law that the number of parameters in an efficient LLM scales as $D^{\gamma}$ where $D$ is the size of the training data and $\gamma\in[0.44,0.72]$, suggesting the existence of more efficient architectures. Against this backdrop, we propose recurrent transformers, combining the efficacy of transformers with the efficiency of recurrent networks, progressively applying a single transformer layer to a fixed-width sliding window across the input sequence. Recurrent transformers (a) run in linear time in the sequence length, (b) are memory-efficient and amenable to parallel processing in large batches, (c) learn to forget history for language tasks, or accumulate history for long range tasks like copy and selective copy, and (d) are amenable to curriculum training to overcome vanishing gradients. In our experiments, we find that recurrent transformers perform favorably on benchmark tests. In our experiments, we find that a single-layer recurrent transformer can match the performance of multi-layer transformers on benchmark tests, delivering comparable accuracy with fewer parameters and lower computational cost.

### v5

@@ 5 Recurrent Transformers
With our theoretical results as backdrop, we propose an architecture that converts any transformer into a recurrent network in order to combine the efficacy of the attention mechanism in the transformer, with the efficiency of recurrent networks. Specifically, let $X=(x_{1},...x_{i},...x_{t})$ be an input sequence, where each $x_{i}$ is a block of vectors embedding the corresponding sequence of non-overlapping and contiguous input tokens $w_{i,1},w_{i,2},...w_{i,k}$, with $k\geq 1$ being the block size. Let $\tau$ denote the function computed by a regular transformer, e.g., VSPUJGK17, in that on input sequence $X$ the transformer outputs $\tau(X)=(y_{1},y_{2},...y_{t})$. We assume that the $x_{i}$ and $y_{i}$ consist of vectors of the same dimension, and that $\tau$ preserves causality in that $y_{i}$ does not depend on $x_{i+1},x_{i+2}...x_{t}$. The recurrent transformer based on $\tau$ is as follows, where the comma between two sequences concatenates them:
@@ 5.1 Computational Complexity
Let $r,d$ and $l$ be the number of layers, the embedding dimension, and the sequence length respectively of a regular transformer and $k$ the block size of a recurrent transformer. We note that the computational complexity of the regular transformer is $r(l^{2}d+ld^{2})$, from which it is easy to see that of the recurrent transformer is $r(4dlk+2ld^{2})$.
@@ 6 Experimental Results
We now test the performance of recurrent transformers on several datasets. The goal of our experiments is to test whether smaller single-layer recurrent transformers can match the performance of larger multi-layer transformers at lower computational cost. Owing to budge constraints, all of our experiments were run on an M4 Mac Mini with 16GB memory. All of our experiments use positional encoding of tokens and the AdamW optimizer.
@@ 6.1 Long Range Image Classification
We compare the performance of the recurrent transformer against regular transformers on the “Long Range Arena” image classification dataset of tay2020long:
“This task is an image classification task, where the inputs are sequences of pixels. In other words, an $N\times N$ image is flattened to a sequence of length $N^{2}$ pixels. …To simplify the setup, we map the input images to a single gray-scale channel where each pixel is represented with an 8-bit pixel intensity (vocabulary size of 256). In LRA, we use the CIFAR-10 dataset…”
Specifically, the CIFAR-10 dataset, krizhevsky2009learning, comprises a training set of 50,000 images of $32\times 32$ pixels, and 10,000 test images across 10 categories. For this task, the images are converted into 256 grayscale values and fed as an input sequence of 1024 tokens across a vocabulary of 256. The output is a single token across a vocabulary of 10 items representing the classification label of the images. Table 1 specifies the base transformer layer in our comparison.
@@
Regular Transformer 4 40 9.1E4 1.38E8 1.68 (0.02) 41.2 (0.7)
Regular Transformer 1 40 5.3E4 3.46E7 1.75 (0.02) 39.3 (0.8)
Recurrent Transformer 1 120 5.3E4 6.29E6 1.50 (0.01) 47.3 (0.4)
@@
Table 2 shows the results of our experiments for this dataset, comparing a 4-layer and 1-layer regular transformer against a recurrent transformer with block size of $32$. The "Ops/sample" column reflects our computational complexity estimates of section 5.1. The Validation Loss reported is the average across 100 samples and in parentheses, the corresponding standard error, e.g., 1.5 (0.01) implies that with 95% confidence, the validation loss is within $\pm$ 0.02 of 1.5. All three models were run for 600,000 samples, and the recurrent transformer could handle larger batch sizes than the 4-layer transformer in the available memory. The recurrent transformer outperforms on loss and accuracy, requiring less than $5\%$ of the compute of the 4-layer regular transformer. Of interest, the parameter $\alpha$ of Equation 6 converges rapidly to $\approx 0.45$ during training, suggesting accumulation of history aids in solving this task. Figure 1 shows the loss and accuracy during training.
@@ 6.2 Copy and Selective Copy with Curriculum Training
We test the performance of the recurrent transformer on the copy and selective copy tasks of gu2023mamba. The copy task involves 16 tokens, followed by 4096 noise tokens, followed by 16 tokens to trigger recall of the first 16 tokens, for a total sequence of 4128 tokens. The tokens are drawn from an alphabet of 16 unique characters, of which 2 characters are reserved for the noise and recall tokens respectively. Successful copy would reproduce the first 16 tokens in the original order against the last 16 recall tokens. The selective copy task randomly intersperses the 16 tokens to be copied amongst 4096 noise tokens, followed by 16 tokens to trigger recall of the 16 interspersed tokens, for a total sequence of 4128 tokens.. Successful copy would reproduce the 16 interspersed tokens in their original order against the last 16 recall tokens.
It is known that copy tasks are easy for transformers, but challenging for recurrent networks, requiring $\approx 100M$ parameters, e.g., gu2023mamba, ren2024exploring. In our experiments, the surprisingly small recurrent transformer of Table 3 with 0.5M parameters solves the tasks, with block size $32$.
@@
Training the recurrent transformer on the copy tasks is challenging due to vanishing gradients, e.g., hochreiter1991untersuchungen. We overcome this challenge with curriculum training, e.g., sodhani2020toward. Specifically, the training process begins with say, 128 noise tokens, trains until the validation loss drops down to say 0.3, then doubles the number of noise tokens. This process is repeated until the number of noise tokens reaches 4096, at which point the training process drives the validation loss down to the desired minimal level for completion. During training, the learning rate decays as a function of the sequence length and the validation loss.
Table 4 shows the results of our experiments. The Ops/sample reported in the table is when each sample has 4096 noise tokens, which is the case during inference. The Validation Loss and Validation Accuracy are averaged over 100 samples and the numbers in parentheses are the respective standard error. For comparison, the estimated Ops/sample is $\approx 1\%$ of a regular transformer with four layers of Table 3. The values of the accumulation parameter $\alpha$ in the table suggest a balance between forget and accumulate to solve this task. Figure 2 shows the loss and accuracy for the selective copy task during the training process. The spikes in the figures are the result of the shocks introduced by doubling the number of noise tokens during curriculum training. The validation loss and accuracy are better than the corresponding training loss and accuracy due to the dropout during training.
@@ 6.3 Natural Language Processing
We compare the performance of the recurrent transformer against regular transformers on the nanoGPT Shakespeare model and dataset of karpathy2023. Table 5 specifies the base transformer layer in our comparison. The recurrent transformer has a block size $16$. Table 6 shows the results of our experiments comparing the performance of a 1-layer and 6-layer regular transformer operating on the full sequence length, against a 1-layer recurrent transformer. The Validation Loss is averaged over 40 samples and the number in parentheses is the respective standard error. The Ops/sample entries reflect the computational complexity estimates of Section 5.1. Figure 3 shows the Training Loss and Validation Loss for our experiments against the total training cost per Section 5.1. Although a fifth of the size, the recurrent transformer matches the performance of the multi-layer regular transformer in this experiment, requiring more samples to train but at a lower cost per sample, and therefore about the same total training cost. However, the recurrent transformer requires only about $20\%$ the inference cost of the multi-layer regular transformer. Of interest, the parameter $\alpha$ of Equation 6 converges rapidly to zero during training. Owing to budge constraints, we are unable to test larger models.
@@
Regular Transformer 6 64 11E6 3.8E8 1.47 (0.01)
Regular Transformer 1 64 1.E6 6.3E7 1.58 (0.01)
Recurrent Transformer 1 64 1.9E6 8.2E7 1.47 (0.01)
@@ 7 Summary
Recent LLMs have hundreds of billions of parameters consuming vast resources. Furthermore, the so-called "AI scaling law" for transformers suggests that the number of parameters must scale linearly with the size of the data. In response, we inquire into efficient LLMs, i.e. those with the fewest parameters that achieve the desired accuracy on a training corpus. Specifically, by comparing theoretical and empirical estimates of the Kullback-Leibler divergence, we derive a natural AI scaling law that the number of parameters in an efficient LLM scales as $D^{\gamma}$ where $D$ is the size of the training data and $\gamma\in[0.44,0.72]$, suggesting the existence of more efficient architectures. Against this backdrop, we propose recurrent transformers, combining the efficacy of transformers with the efficiency of recurrent networks, progressively applying a single transformer layer to a fixed-width sliding window across the input sequence. Recurrent transformers (a) run in linear time in the sequence length, (b) are memory-efficient and amenable to parallel processing in large batches, (c) learn to forget history for language tasks, or accumulate history for long range tasks like copy and selective copy, and (d) are amenable to curriculum training to overcome vanishing gradients. In our experiments, we find that recurrent transformers perform favorably on benchmark tests. In our experiments, we find that a single-layer recurrent transformer can match the performance of multi-layer transformers on benchmark tests, delivering comparable accuracy with fewer parameters and lower computational cost.
8 Reproducibility
All code available on github.